On January 1, 2019, Tower company issues $15,000 in bonds for $14,700. They were 6 year bonds with a coupon rate of 9% and pay semiannual interest. The straight-line method is used to amortize the bond discount. On June 30, 2019 when the first payment is made, how much interest Expense will be recorded?
Answer #1

Answer : $700

Interest Expenses = Cash interest + Discount Amortized

Cash interest = 15000*9%*6/12 = 675

Bond Discount Amortization = (Face Value - Issue Price )/ Number of Payments

Bonds Discount Amortized with Each Payment = (15000-14700)/12 = 25

Interest Expenses with First Payment = 675+25 = 700 (Answer)

I